# AgMMU: A Comprehensive Agricultural Multimodal Understanding and Reasoning Benchmark

### v1

@@ 3.1 Dataset Curation
Then we employ GPT [1] to convert these facts into MCQs and OEQs. Depending on the knowledge types of the fact, we give the LLM a unique prompt to construct a corresponding question. We prompt it to generate three wrong answers and use the extracted fact as the ground truth to form one correct answer. To ensure fairness of the question, we extract “background information” with LLaMA-405B from the original AskExtension question which contains relevant information that the user provided. We also include the month,and location information in the question. This process gives us 5,460 unique MCQs and OEQs paired with 12,481 unique images.
Ethical Compliance.   We use automatic procedures assisted with human verification to erase the information that could leak the name, gender, address, or any other personal information for both users and experts. We will only distribute the processed AgMMU instead of the source data from the Internet to protect the privacy of the users and experts. To the best of our verification, we have not found any human faces in the benchmark images. Since we only keep the questions related to agriculture, our benchmark does not contain gender, race, or other types of societal biases.
3.2 Specialty of AgMMU
@@
Distribution and Coverage.   In Figure 3, we show the distribution of agriculture sub-domains and knowledge types in AgMMU. We have two key observations: (1) The original dataset we obtained from the Internet is severely skewed in sub-domains: Sub-domains like nutrient deficiency and generic identification domains have significantly less samples than dominating domains such as disease advice or growing advice. This imbalance is inevitably carried over to knowledge types after the knowledge extraction stage. (2) However, in AgMMU, we demonstrates a much more balanced distribution with our dataset curation pipeline. This is exceptionally important for our evaluation set, where we aim for proposing an all-encompassing benchmark that examines different faces of models without inductive bias.
@@
Realistic Images.   Our AgMMU also features significant challenges in the quality and number of images. The images in our benchmark are generally uploaded by users in various quality, resolution, and aspect ratios. This differs from the photography-level images in datasets like iNaturalist [18]. As shown in Figure 1, AgMMU realistically reflect the multimodal everyday scenarios encountered by gardeners and farmers. Our benchmark also exhibits the necessity of multi-image understanding challenge, since the user uploaded several images and the experts rely on a joint reasoning of them to make the final conclusion.
@@
Disease/Issue Id.
Insect/Pest Id.
Species Recog.
Management Instr.
Symptom/Visual Descr.
@@
Disease/Issue Id.
Insect/Pest Id.
Species Recog.
Management Instr.
Symptom/Visual Descr.
@@
GPT-4o [1] 21.02 26.13 19.79 10.37 11.65 17.79 80.51 84.46 86.23 90.42 84.64 85.25
Gemini 1.5 Pro [33] 25.17 29.77 24.88 14.15 15.91 21.98 76.25 79.92 81.07 88.08 76.94 80.42
Claude 3 Haiku [40] 10.51 18.09 20.34 1.63 15.73 13.26 62.14 68.54 49.89 81.76 52.03 62.00
@@
Cambrian-8B [46] 6.78 17.29 18.55 1.96 5.00 9.92 63.66 60.84 57.49 73.28 76.38 65.81
InternVL2-8B [43] 4.20 15.61 8.13 1.79 6.64 7.27 50.95 58.14 58.34 73.13 61.31 60.17
LLaMA-3.2 [10] 11.56 23.96 21.25 4.02 18.77 15.91 65.07 69.29 75.21 78.50 78.08 73.32
LLaVA-NeXT-8B [25] 4.81 15.98 20.56 0.73 1.86 8.79 62.18 67.18 60.03 75.31 71.19 66.71
LLaVA-OneVision-8B [21] 5.97 16.02 15.39 2.59 8.90 9.77 69.07 70.93 67.62 81.87 75.02 72.53
LLaVA-1.5-7B [24] 3.98 14.64 23.07 0.97 9.45 10.42 62.66 61.24 57.18 70.64 71.32 64.16
LLaVA-1.5-13B [24] 5.19 15.70 24.21 2.58 10.72 11.68 66.57 58.73 60.41 78.29 72.26 66.73
Qwen-VL-7B [4] 4.97 17.94 22.34 1.89 8.24 11.08 56.24 58.31 60.65 69.64 67.70 62.34
VILA1.5-13B [23] 10.32 21.12 23.58 2.98 11.02 13.80 60.18 62.76 60.71 70.78 61.63 63.00
@@
4 Experiments
We perform comprehensive evaluations on diverse vision-language models (VLMs), including both proprietary and open-sourced models. Our evaluation primarily focuses on zero-shot performance, utilizing either publicly available APIs or author-provided checkpoints to reflect the inherent capabilities of each model without task-specific training. Furthermore, we present fine-tuning results on LLaVA-1.5 to demonstrate the importance of our knowledge base AgBase-200K for optimizing performance on agriculture tasks. All experiments are conducted using NVIDIA A6000 GPUs, and further details regarding model configurations are available in the supplementary material.
4.1 Baselines and Metrics
Baselines.   We include a broad range of state-of-the-art VLMs to ensure robust comparisons. For open-sourced models, we prioritize architectures with comparable parameter numbers for a fair and meaningful comparison. Moreover, we include the most recent, largest, and highest-performing checkpoints up to the date of our evaluation questions, including LLaMA-3.2 [10], LLaVA-1.5 [24], LLaVA-NeXT [25], LLaVA-OneVision [21], Cambrian-1 [46], InternVL-2 [43], Qwen-VL [4], and VILA [23], as well as proprietary models such as GPT-4o [1], Claude-3.5 [40], and Gemini1.5-Pro [41]. Detailed configurations and model settings are in the supplementary material.
Evaluation.   We evaluate the models on multiple-choice questions (MCQs) and open-ended questions (OEQs). To ensure a robust and fair evaluation, we have developed a systematic approach to minimize potential variations in model responses due to intermediate generations, capitalization differences, or formatting inconsistencies.
For MCQs, we report accuracy as the primary evaluation metric for MCQs within the benchmark, reflecting the models’ ability to perform choice tasks relevant to agricultural scenarios. We score model response by conducting basic string pre-processing and matching the model letter prediction with the true letter associated with the ground truth.
For OEQs, we implement the LLM-as-judge methodology using Qwen-7B to grade the answers. The grading is divided into two categories: (1) few-word responses and (2) multi-statement responses. Few-word responses consist of the categories pest identification, disease identification, and species identification. These responses consist of two to three words and are evaluated directly against the correct answer. We use a structured scoring rubric following SimpleQA [50] and assign the grades of “correct”, “incorrect”, “partially correct”, and “not attempted”. We use the harmonic mean to calculate the final grade. Multi-statement responses, on the other hand, consist of the categories management instructions, and symptom descriptions. These categories cannot simply be evaluated directly against the correct answer because they usually consist of multiple unique and standalone statements. On average, the type of management instruction of responses has 2.73 correct unique statements and symptom/visual description type responses has 2.74 statements. Since we find the complexity of separating answers into unique statements and grading them individually to be too complex for a single inference of our LLM judge, we split the process into two steps. (1) In the first step, we ask the LLM judge to split the answer into standalone statements. And (2) in the next step, we have the LLM judge compare every statement in the correct answer with every statement in the predicted response using a similar grading rubric, which assigns the grades “correct”, “incorrect”, “partially correct”, and “irrelevant”.
The highest grade assigned to a predicted statement is saved as its score. For each question, we normalize the correct and partially correct predicted statements with the total correct statements. Then we use the harmonic mean to calculate the final grade [50].
@@ 4.2 Zero-shot Evaluation Results
Very Challenging for Existing Models  : The AgMMU benchmark proves to be notably challenging across all evaluated models, and even the most advanced systems achieve moderate performance levels. This demonstrates that AgMMU reflects the inherent complexity of visual understanding and reasoning tasks in agriculture, where models must simultaneously process intricate visual details and apply domain-specific knowledge.
Performance Across Different Question Types  : For MCQs, the most challenging knowledge type categories are consistently those involving insect and disease identification, or species recognition. Management instruction and visual description questions present slightly smaller challenges, with most VLMs’ scores often 5-10% above the more complex recognition tasks. For OEQs, the management instruction and visual description types of questions pose more challenges, due to their long-response nature and the need for more reasoning. However, there is still significant variation between models as to what tasks are the most difficult, suggesting that specialized agricultural knowledge and visual understanding capabilities are not uniformly distributed across model architectures.
Figure 4: Error Analysis. The most common errors made by VLMs are knowledge errors. CoT represents samples that are originally wrong due to false reasoning, but corrected with chain-of-thought prompting. IPI, DII, SVD, SI, and MI are short for five question types as explain in Sec 3.1.
Open-ended Question Performance  : The performance difference between OEQs and MCQs is substantial, with the performance of the OEQ trailing 40 to 50% behind the MCQ results. These results suggest that free-response tasks, which do not allow the elimination process with candidate answers, require a much more nuanced understanding and generation of information and remain a considerable hurdle for open-sourced and closed-source models.
Between Closed- and Open-sourced Models  : We observe a distinct performance gap between closed-source and open-source models, where proprietary models generally demonstrate superior performance. Among all models, Gemini 1.5 Pro and GPT-4o lead in all tasks. For open-source models, LLaMA-3.2 is the overall best performer across all subdomains except management instructions where Cambrian-8B is the best. On average, open-sourced models trail behind their closed-source counterparts by more than 10-20%.
Figure 5: Error Demonstration. We show two common evaluation errors, including the lack of knowledge to answer the question (left), and the wrong perception of the image (right).
@@ 4.3 Error Analysis
We analyze the types of error that occurred during the VLM evaluation. We select 150 wrong evaluation samples and ask the VLM to provide the chain-of-thought reasoning before providing the final answer, and judge the error types with human evaluators. Figure 4 demonstrates quantitative results, and Figure 5 shows two common error types.
Knowledge Error (56.5%).   A knowledge error occurs when the VLM does not have the knowledge necessary to reach the correct answer given the context. In these cases, the model may identify relevant visual features and reinforce important context but cannot conclude the correct answer. For example, in management questions, knowledge errors often lead to vague or non-specific responses that fail to engage with the particular nuances of the given situation.
Figure 6: Open-ended questions (OEQ) examples demonstrating the effectiveness of our fine-tuned LLaVA model, which accurately identifies issues that both GPT-4o and the original LLaVA fail to recognize in zero-shot scenarios. In diverse agricultural categories: disease, species, and pest identification, our fine-tuned model exhibits enhanced diagnostic accuracy.
Figure 7: Finetuning on our knowledge base AgBase-200K significantly benefits the agricultural knowledge and factuality of LLaVA, indicating the potential of our training set for developing knowledge-intensive VLMs.
Perceptual Error (18.5%).   A perceptual error occurs when the VLM fails to recognize a primary visual characteristic or gives a directly incorrect description of a visual characteristic. This category highlights the incapability in visual understanding of VLMs.
Chain-of-Thought (22.8%).   The chain-of-thought (CoT) category includes instances where the VLM initially provides an incorrect response, but corrects itself when asked to explain its reasoning. The amount of samples in this category emphasizes the importance of reasoning procedures in obtaining the correct answers for many questions on the AgMMU benchmark.
Reject to Answer (2.2%).   In rare situations, we find that the VLM refuses to answer the question.
@@ 4.4 Finetuning Evaluation Results
Analysis.   Figure 6 and 7 demonstrate the effectiveness of fine-tuning a VLM using our AgBase-200K on more challenging open-ended questions. Figure 7 shows that the fine-tuning with our knowledge base brings a significant and consistent improvement across all question domains, leading to an overall 4.7% average improvement. This performance boost highlights the effectiveness of our large-scale knowledge base in improving VLMs in connecting facts with images. Compared to Table 1, having such a training set is an important benefit of our dataset. With the large-scale database, AgMMU is the first robust dataset for such domain-specific fine-tuning. The substantial performance gains observed emphasize the critical role of the dataset in improving the understanding of the model of complex domain-specific visual input and reasoning tasks.
@@ 5 Conclusion
In this work, we introduced AgMMU, the first benchmark for evaluating vision-language models (VLMs) in agriculture, a field that requires precise visual interpretation and expert knowledge. Our dataset spans core agricultural tasks, including symptom recognition, species and pest identification, and management instructions, and is created from more than 116,000 real-world expert-user interactions. Through a three-step curation framework, we build a high-quality AgMMU evaluation set, demonstrating a great challenge for current VLMs. To support model development, we introduce AgBase-200K, a training set of 205,399 knowledge entries aimed at improving model accuracy. We hope that the combination of AgMMU and AgBase can support the community in evaluating and developing stronger knowledge-intensive VLMs.
Acknowledgments.
We thank Kastan Day and Rohan Marwaha for the fruitful discussion. We thank the AskExtension team for helping us with the data curation. This work was supported in part by the AIFARMS National AI Institute, USDA-NIFA Award 2020-67021-32799, NSF Grant 2106825, the University of Illinois Discovery Partners Institute, the Center for Digital Agriculture at the University of Illinois, the Amazon-Illinois Center on AI for Interactive Conversational Experiences, and the IBM-Illinois Discovery Accelerator Institute. This work used computational resources, including the NCSA Delta and DeltaAI supercomputers through allocations CIS230012 and CIS240133 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, as well as the TACC Frontera supercomputer, Amazon Web Services (AWS), and OpenAI API through the National Artificial Intelligence Research Resource (NAIRR) Pilot.
Supplementary Material
[…]

### v2

@@ 3.3 Additional Properties of AgMMU
Distribution and Coverage. In Fig. 3, we show the distribution of agricultural subdomains and knowledge types in AgMMU. We have two key observations: (1) The original conversations we obtained are severely skewed in subdomains: Subdomains like nutrient deficiency and generic identification domains have significantly fewer samples than dominating domains such as disease advice or growing advice. This imbalance is inevitably carried over to knowledge types after the knowledge extraction stage. (2) However, in AgMMU, we demonstrate a much more balanced distribution with our dataset curation pipeline. This is exceptionally important for our evaluation set, where our aim is to propose an all-encompassing benchmark that examines different aspects of models without inductive bias.
@@
Realistic Images.   Our AgMMU also features significant challenges in the quality and number of images. The images in our benchmark are generally uploaded by users in various qualities, resolutions, and aspect ratios. This differs from the photography-level images in datasets like iNaturalist [19] or the textbook or web-document figures like manually curated CROP [54] and CDDM [28]. As shown in Fig. 1, AgMMU realistically reflect the multimodal everyday scenarios encountered by gardeners and farmers. Our benchmark also exhibits the necessity of multi-image understanding challenge, since the user uploaded several images and the experts rely on a joint reasoning of them to make the final conclusion.
4 Experiments
We perform comprehensive evaluations on diverse vision-language models (VLMs), including both proprietary (closed-source) and open-sourced models. Our evaluation primarily focuses on zero-shot performance, utilizing either publicly available APIs or author-provided checkpoints to reflect the inherent capabilities of each model without task-specific training. Furthermore, we present fine-tuning results on LLaVA-1.5 to demonstrate the importance of our development set AgBase for optimizing performance on agriculture tasks, and highlight some unique observations. All experiments are conducted using NVIDIA A6000 GPUs, and further details regarding model configurations, API choices, and implementation details are available in the supplementary material.
4.1 Baselines
We include a broad range of state-of-the-art VLMs to ensure robust comparisons. For open-sourced models, we prioritize architectures with comparable parameter numbers for a fair and meaningful comparison. Moreover, we include the most recent, largest, and highest-performing checkpoints up to the date of our evaluation questions, including LLaMA-3.2 [11], LLaVA-1.5 [25], LLaVA-NeXT [26], LLaVA-OneVision [22], Cambrian-1 [47], InternVL-2 [44], Qwen-VL [4], and VILA [24], as well as proprietary models such as GPT-o4-mini [1], Claude-3.5 [41], and Gemini1.5-Pro [42]. Detailed configurations and model settings are provided in the supplementary material.
4.2 Evaluation Metrics
We evaluate the models on multiple-choice questions (MCQs) and open-ended questions (OEQs). To ensure a robust and fair evaluation, we have developed a systematic approach to minimize potential variations in model responses due to intermediate generations or formatting inconsistencies.
For MCQs, we report accuracy as the primary evaluation metric. We score model response with string pre-processing and matching the predicted letter with the ground truth letter.
For OEQs, we implement the LLM-as-judge methodology using GPT-4.1 to grade the answers. The prompts and style are initialized from SimpleQA [50] setting, but adapted for AgMMU scenarios. The grading is divided into two categories: (1) short-form responses and (2) multi-statement responses. The short-form responses correspond to the questions from pest identification, disease identification, and species identification, normally containing several words, and are evaluated directly against the correct answer. We assign the grades of “correct”, “incorrect”, “partially correct”, and “not attempted,” then use the harmonic mean to calculate the final grade of $\text{Correct}/\text{Total}Correct / Total$” and “Correct/(TotalPartially Correct)CorrectTotalPartially $\text{Correct}$/($\text${Total}-$\text${Partially Correct})Correct / ( Total - Partially Correct )”, similar to SimpleQA. The multi-statement responses, on the other hand, are for management instructions and symptom descriptions questions. These categories usually consist of multiple unique and standalone statements: On average, management instruction responses contain 2.7 factual statements, and symptom/visual description responses contain 1.8 factual statements. Therefore, we have our LLM judge (a) divide both predicted and ground-truth responses into individual statements; and (b) grade each of them as a short-form response; then (c) normalize the grades according to the number of statements per question to calculate the final harmonic mean score as short-form responses.
@@ 4.2 Evaluation Metrics
Disease
Insect/Pest
Species
Management
Symptom
@@
Disease
Insect/Pest
Species
Management
Symptom
@@
GPT-o4-mini [1] 47.7 48.8 69.6 16.2 2.6 37.0 77.9 85.4 90.3 93.8 84.3 86.5
Gemini 1.5 Pro [34] 37.6 51.7 69.9 32.5 16.9 41.7 76.2 81.1 82.8 88.1 76.9 82.4
Claude 3 Haiku [41] 25.1 20.6 32.2 1.0 13.2 18.4 62.1 71.2 52.8 81.5 52.0 63.8
@@
LLaVA-1.5-7B [25] 4.9 22.3 30.6 1.4 5.0 12.8 61.9 64.6 67.6 77.3 71.5 69.0
LLaVA-NeXT-8B [26] 6.4 27.2 31.1 2.6 0.3 13.5 61.9 70.1 57.9 78.5 67.6 67.5
LLaVA-OneVision-8B [22] 9.1 24.1 34.1 5.4 6.2 15.8 65.7 72.9 71.2 85.9 78.8 75.4
LLaVA-1.5-13B [25] 8.2 22.1 31.8 1.1 5.9 13.8 64.7 67.4 65.5 80.4 73.7 70.8
Cambrian-8B [47] 7.9 30.9 27.9 2.2 1.9 14.2 65.0 70.1 59.3 79.1 86.0 72.8
InternVL2-8B [44] 8.5 27.0 22.1 1.0 1.9 12.1 55.5 61.0 60.4 74.7 64.3 63.5
Qwen-VL-7B [4] 10.2 25.2 36.1 4.1 6.5 16.4 55.5 61.0 60.4 74.7 64.3 63.5
VILA1.5-13B [24] 11.0 34.1 36.8 4.9 6.5 18.7 61.8 63.2 60.1 73.9 59.8 63.7
LLaMA-3.2-11B [11] 22.9 32.5 46.2 5.7 17.1 24.9 66.2 75.0 78.6 89.6 79.9 78.3
@@ 4.3 Zero-shot Evaluation Results
Very Challenging for Existing Models.   AgMMU proves to be a very challenging benchmark across all models evaluated, and even the most advanced systems achieve moderate performance levels. Such challenges are notable, especially on open-ended OEQs, where models have to recall the knowledge precisely without relying on any pre-input options. These observations also reveal the necessity of improving the agricultural expertise of VLMs and mixing agricultural data for VLM training, where our AgBase could be helpful.
Performance Across Question Types.   On MCQs, the models can rely on options, so it is significantly easier than OEQs. The level of >>>70% accuracy is also on par with previous agricultural benchmarks [54, 28] in MCQs. Therefore, we primarily advocate for the more challenging OEQ setting, which is also more aligned with a VLM in the real world. As observed in Table 2, the management instruction and symptom/visual description types of questions pose more challenges, due to their multi-statement and long-response nature and the need for more reasoning. However, there is significant variation between models as to what tasks are the most difficult, suggesting that specialized agricultural knowledge and visual understanding capabilities are not uniformly distributed across model architectures.
(a) Error Analysis.
(b) Error Demonstration.
Figure 4: (a) The most common errors made by VLMs are knowledge errors. CoT represents samples that are originally wrong due to false reasoning, but corrected with chain-of-thought prompting. IPI, DII, SVD, SI, and MI are short for five question types as explained in Sec 3.2. (b) We show two common evaluation errors, including the lack of knowledge to answer the question (top), and the wrong perception of the image (bottom).
Closed- and Open-sourced Models.   We observe a distinct performance gap between closed-source and open-sourced models, where closed-source models generally demonstrate superior performance. Among all models, Gemini 1.5 Pro and GPT-o4-mini lead in all tasks. For open-sourced models, LLaMA-3.2 is the overall best performer across all subdomains. On average, open-sourced models trail behind their closed-source counterparts by more than 10 to 20%, likely due to the lack of substantial agricultural data during the large-scale training phases.
@@ 4.4 Error Analysis
To achieve an in-depth understanding of the internal procedures of VLMs in answering AgMMU questions, we analyze the types of error that occurred during the evaluation. We select 171 wrong evaluation samples and ask the VLMs to provide the chain-of-thought (CoT) reasoning before providing the final answer, and judge the error types with human evaluators. Fig. 4(a) demonstrates quantitative results, and Fig. 4(b) shows the qualitative results of the most common knowledge and perceptual errors.
Perceptual Error.   A perceptual error occurs when the VLM fails to recognize a primary visual characteristic or gives a direct incorrect description of a visual characteristic as in Fig. 4(b). This category highlights limitations in the visual understanding capabilities of VLMs. Perceptual errors constitute 34.3% (Gemini), 37.5% (LLaMA), and 22.2% (GPT-o4) of total errors.
Knowledge Error.   A knowledge error occurs when the VLM does not have the essential knowledge to reach the correct answer. In these cases, the model may identify relevant visual features and reinforce it, but without reaching the correct answer. For example, in management questions, the model might give vague responses that do not engage with the particular nuances of the given situation (Fig. 4(b)). Knowledge errors constitute the highest proportion of errors for all Gemini (48.6%), LLaMA (42.9%), and GPT-o4 (51.1%).
Chain-of-Thought (CoT).   This category includes instances where the VLM initially provides an incorrect response but corrects itself after being asked to explain its reasoning. The proportion of CoT errors underscores the importance of reasoning processes for accurate performance on our benchmark. Specifically, CoT accounts for 15.7% (Gemini), 16.1% (LLaMA), and 26.7% (GPT-o4) of total errors, indicating that these models benefit significantly from explicit reasoning steps.
Reject to Answer.   Occasionally, VLMs refuse to answer a question entirely, reflecting uncertainty or lack of confidence in their predictions. This occurs rarely, constituting 1.4% (Gemini) and 3.6% (LLaMA), with no instances observed in GPT-o4.
The differing error patterns across Gemini, LLaMA, and GPT-o4 reveal distinct model behaviors. Perceptual ability forms the foundation: models like LLaMA, with weaker visual grounding, show more perceptual errors—often hallucinating features to match textual cues. In contrast, Gemini accurately identifies main visual elements without hallucination but frequently makes knowledge errors, indicating gaps in domain understanding despite solid perception. GPT-o4, with strong perception and knowledge, rarely makes blatant errors, but its reasoning steps can still misfire.
Figure 5: The effectiveness of AgBase fine-tuning on OEQ examples. After simple fine-tuning, the LLaVA model can accurately identify issues that GPT and the original model fail to recognize in zero-shot scenarios.
@@ 4.5 Finetuning Evaluation Results
Figure 6: Fine-tuning on AgBase boosts the agricultural knowledge understanding for LLaVA-1.5, indicating the potential of our development set. IPI, DII, SVD, SI, and MI are the short for the five question types as explained in Sec. 3.2.
@@
Analysis.   Fig. 5 and Fig. 6 demonstrate the effectiveness of fine-tuning VLMs using our AgBase on the more challenging open-ended questions. Fig. 6 shows that fine-tuning with our knowledge base significantly improves the model’s capability of understanding the images and correctly responding with agriculture knowledge, leading to up to 11.6% improvement on average. This performance boost highlights the effectiveness of our large-scale knowledge base in improving VLMs and the necessity of collecting agricultural-related data for future VLMs. Meanwhile, the smaller improvement on management instruction and symptom description questions indicates the complexity of multi-statement questions, motivating the need for more data collection and better model training than simple fine-tuning.
@@ 5 Conclusion
In this work, we introduced AgMMU, the first benchmark for evaluating vision-language models (VLMs) in agriculture, a field that requires precise visual interpretation and expert knowledge. Our dataset spans core agricultural tasks, including symptom recognition, species and pest identification, and management instructions, and is created from more than 116,231 real-world expert-user interactions. Through a three-step curation framework, we build a high-quality AgMMU evaluation set, demonstrating a great challenge for current VLMs. To support model development, we introduce AgBase, a training set of 57,079 knowledge entries aimed at improving model accuracy. We hope that the combination of AgMMU and AgBase can support the community in evaluating and developing stronger knowledge-intensive VLMs.